# Self-Modeling Robots by Photographing

### v1

@@ Abstract
Self-modeling enables robots to build task-agnostic models of their morphology and kinematics based on data that can be automatically collected, with minimal human intervention and prior information, thereby enhancing machine intelligence. Recent research has highlighted the potential of data-driven technology in modeling the morphology and kinematics of robots. However, existing self-modeling methods suffer from either low modeling quality or excessive data acquisition costs. Beyond morphology and kinematics, texture is also a crucial component of robots, which is challenging to model and remains unexplored. In this work, a high-quality, texture-aware, and link-level method is proposed for robot self-modeling. We utilize three-dimensional (3D) Gaussians to represent the static morphology and texture of robots, and cluster the 3D Gaussians to construct neural ellipsoid bones, whose deformations are controlled by the transformation matrices generated by a kinematic neural network. The 3D Gaussians and kinematic neural network are trained using data pairs composed of joint angles, camera parameters and multi-view images without depth information. By feeding the kinematic neural network with joint angles, we can utilize the well-trained model to describe the corresponding morphology, kinematics and texture of robots at the link level, and render robot images from different perspectives with the aid of 3D Gaussian splatting. Furthermore, we demonstrate that the established model can be exploited to perform downstream tasks such as motion planning and inverse kinematics.
@@ 1 Introduction
High-quality models are essential for effective robot planning, control, and visualization. Traditional modeling approaches demand substantial human effort to gather the necessary parameter information of the robot, and over time, the quality of these models tends to deteriorate due to various factors, such as unexpected damage, joint malfunctions, and wear. Therefore, the development of methods that enable robots to collect data and learn their own models autonomously is critical for ensuring the successful execution of tasks and advancing the level of machine intelligence. Inspired by human behavior (Rochat, 2003), the concept of robot self-modeling has recently gained attention among researchers (Chen et al., 2022; Díaz Ledezma and Haddadin, 2023). A robot self-model is a general-purpose representation of a robot’s physical shape and structure that can be acquired at any time without human intervention (Kwiatkowski, 2022). Through self-modeling, robots can autonomously and continuously update their knowledge of both morphology and kinematics, which not only ensures that the model information remains up to date and adaptable to various unforeseen circumstances but also simplifies the modeling process for newly designed robots.
@@
Early approaches focused on designing task-specific self-modeling techniques, (Kwiatkowski and Lipson, 2019; Chen et al., 2021; Hang et al., 2021) such as modeling end-effector positions (Mathew et al., 2014), tilt angles (Bongard et al., 2006), or joint velocities (Sanchez-Gonzalez et al., 2018) according to task requirements. However, due to the lack of generality and adaptability in these methods, recent work has shifted toward modeling the whole robot to construct task-agnostic self-models. For instance, Chen et al. (2022) employed visual data with depth information to enable self-modeling of the overall morphology of a 4-degree-of-freedom robot, which was later validated on robots with higher degrees of freedom by Yang et al. (2024a). Their approach allows the robot to determine its spatial occupancy based on input joint configurations but cannot directly compute the kinematic information (e.g., position) of specific links. To address link-level modeling,  Yang et al. (2024a) developed multiple independent models for each robot link and integrated them into a comprehensive self-model. However, these methods require depth information to create self-models, which necessitates expensive devices (e.g., depth cameras and LiDAR) and is hard to obtain in practice. In addition to visual data, inertial measurement unit (IMU) data can also be utilized for the self-modeling of robot (Díaz Ledezma and Haddadin, 2023), which necessitates the attachment of numerous IMU sensors to the robot links. Notably, despite the considerable advancements in robot self-modeling concerning morphology and kinematics, the aforementioned approaches have neglected the modeling of texture features (such as surface color and patterns), which are crucial for the visualization of robots in simulation environments and digital twin systems. Thus, beyond morphology and kinematics, texture modeling remains a significant challenge in achieving comprehensive robot self-modeling.
@@
Given the challenges and high costs of data acquisition in the previously discussed self-modeling methods and the importance of texture modeling, a visual self-modeling approach utilizing only two-dimensional (2D) images presents a more efficient and cost-effective alternative. In this paper, we develop a scheme based on 3DGS (Kerbl et al., 2023) and neural networks for modeling the robot’s morphology, kinematics, and texture. 3DGS is a newly developed technology for 3D scene reconstruction, which represents 3D scenes with a large number of optimizable, flexible, and expressive 3D Gaussian ellipsoids, enabling real-time rendering of scene images. With the assistance of 3DGS, we can acquire 3D information, which is necessary for robot self-modeling from 2D images. An overview of our method is demonstrated in Figure1a. To optimize the 3D Gaussians and kinematic network, the robot only needs to gather multi-view RGB images under various configurations, along with corresponding joint configuration data and camera parameters. The static morphology and texture information of the robot’s links are embedded in the 3D Gaussians, each of which is an ellipsoid defined by its position, covariance matrix, color, and opacity. By clustering these 3D Gaussians, the bones of the robot model are constructed, with transformations of the bones determined by the kinematic network. By feeding the kinematic neural network with joint angles, we can utilize the well-trained self-model to describe the corresponding morphology, kinematics, and texture of robots at the link level and render robot images from different perspectives with the aid of 3DGS. Besides, the self-model can be employed to directly control the robot to perform downstream tasks such as motion planning and inverse kinematics. As illustrated in Figure1b, compared to previous methods, the proposed method relies solely on easily accessible data (e.g., multi-view RGB images) for learning while producing a high-quality, texture-aware, and link-level self-model of robots.
@@
Figure 1: Robot self-modeling from images. a Given a robot whose model is unknown, our method can build a self-model of the robot from multiple RGB images. We capture multi-view photos of the robot at different joint configurations, and train the 3D Gaussians and the kinematic network, which are responsible for representing the static morphology and texture of robots and neural bone deformation, respectively. The well-trained model can be utilized to render robot images with the aid of 3D Gaussian splatting and perform downstream tasks. b Compared to the methods proposed by Chen et al. (2022) and Schulze and Lipson (2024), our method requires no depth information and can establish a high-quality, texture-aware and link-level self-model.
Overall, the contributions of this paper are:
  • We propose a robot self-modeling method that can be built from images without 3D information based on 3DGS. The robot self-model can correctly reconstruct the robot’s morphology, kinematics, and texture.
  • We model the robot at the link level. The self-model automatically identifies different links of the robot and can predict their movement separately.
@@
  • We use the self-model to control the robot. Experiments show that the self-model can control the robot to accomplish downstream tasks without additional training.
@@ 2 Related work
Neural Radiance Field (NeRF) has shown remarkable results in reconstructing scenes from multiple-view photos. Pumarola et al. (2021) proposed the first method to model dynamic scenes with NeRF. To accelerate the training of NeRF, methods such as (Fang et al., 2022; Müller et al., 2022) are proposed. These works motivate researchers to use dynamic NeRF to reconstruct and animate the human body (Noguchi et al., 2021)and animals (Yang et al., 2022). Noguchi et al. (2022) combined NeRF with signed-distance functions (SDFs) to learn both the appearance and the structure of articulated objects, including robots and human bodies, by observing them move from multiple views. They modeled articulated objects with several ellipsoids. They could repose articulated objects by transforming ellipsoids. Uzolas et al. (2024) extracted point clouds of articulated objects from pre-trained NeRF models and used Linear Blend Skinning (LBS) to learn Dynamic NeRF and associated skeletal models jointly. Although these works achieved modeling robots from images or videos, they did not associate robots’ joints with the built morphology models. In other words, they did not model the robots’ kinematics.
@@
Though NeRF shows promising results in 3D reconstruction, it has a significant drawback. To get a high-quality scene reconstruction, it may take a long time. The 3DGS method is designed to solve this problem and can quickly complete training and build high-quality models. Similar to NeRF, researchers extended 3DGS to the reconstruction of dynamic 3D scenes (Wu et al., 2024; Yang et al., 2024c; Liang et al., 2023; Yang et al., 2024b; Lin et al., 2024).  Hu et al. (2024); Qian et al. (2024) proposed to model 3D human avatars with 3DGS. Several other approaches aim to design controllable and editable models with 3DGS. Yu et al. (2024) used pre-defined masks to choose moving parts of the modeled scene. These masks are used to control the model, and only the chosen part can be controlled. Huang et al. (2024) used control points to control the dynamic 3DGS model. Control points are automatically generated during training, which then control the movement of the whole 3DGS model with Linear Blend Skinning. Zhang et al. (2024) proposed to use neural bones to control the 3DGS model, which was similar to the method proposed by Yang et al. (2022). A neural bone is an ellipsoid parameterized by position and radius. When controlling the 3DGS model, neural bones work in the same way as real bones in human and animal bodies.
2.3 Modeling robots with 2D input
@@
Several studies have explored using 2D images for robot self-modeling (Hu et al., 2023; Marques Monteiro et al., 2024). However, these approaches are limited to rendering images of the robot in various states and do not succeed in constructing a comprehensive model of the robot’s morphology. Schulze and Lipson (2024) was the first to solve the robot self-modeling problem using dynamic NeRF and associated the built model with input robot joints. However, NeRF-based self-modeling requires optimizing a dense multi-layer perceptron (MLP) for learning both morphology and kinematics, and uses volumetric ray-marching to render robot images. As a result, NeRF-based self-modeling suffers from problems such as difficulties in model training, high rendering costs, and suboptimal modeling quality. Lou et al. (2024) used a hybrid representation that combines 3DGS and mesh to create digital assets of robots. The model of robots can be controlled by robot joints. However, real Denavit-Hartenberg parameters are required to control the model based on joints, which requires human intervention and fails to model robot kinematics automatically.
@@ 3 Method
Figure 2: An overview of our method. We first train a static 3DGS self-model with images of the robot at zeor pose from various views. Then, we initialize neural ellipsoid bones to model the robot kinematic chain. Using images of the robot at different joint configurations, we train a dynamic self-model. The neural bones are deformed, and the relative 3D Gaussian is controlled by LBS. Finally, the model is rendered into images and compared with the ground truth images.
We propose a novel robot self-modeling method that begins by modeling the robot in a rest pose, which we define as canonical space. The robot is then transformed into various poses, which form the observation space. To implicitly represent the robot’s self-model, we utilize 3D Gaussian splatting (3DGS). The model can be deformed and rendered into images, allowing us to visualize the robot’s morphology from different perspectives. An overview of our method is provided in Figure 2.
@@ 3.1 Preliminaries
3DGS uses 3D Gaussians to represent a static scene (Kerbl et al., 2023). A 3D Gaussian is a small ellipsoid with feature parameters. Typically, a static 3DGS model consists of 10 thousands of 3D Gaussians. Each 3D Gaussian is defined by a full 3D covariance matrix $\Sigma$:
$G(x)\ =\ e^{-\frac{1}{2}{(x)}^{T}\Sigma^{-1}(x)}.$ (1)
The covariance matrix $\Sigma$ can be calculated by the rotation matrix $R$ and scaling matrix $S$ as $\Sigma=RSS^{T}R^{T}$. The rotation matrix is derived from a quaternion $q$. Each Gaussian has an opacity value $\sigma$ and a spherical harmonic coefficient $\xi$ to create view-dependent color. When rendering images, 3D Gaussians are projected to a 2D plane. Given a viewing transformation $W$, the covariance matrix $\Sigma^{\prime}$ in camera coordinates is given by:
@@
where $J$ is the Jacobian of the affine approximation of the projective transformation. The color $C$ of a pixel is computed by alpha blending 3D Gaussians that overlap on the pixel sorted by depth, and the color $C$ is given by:
@@
where $N$ is the number of Gaussians overlapping the pixel, $c_{i}$ is the color of Gaussian $i$ calculated by spherical harmonic coefficient, $\alpha$ is given by evaluating the projected Gaussian with covariance matrix $\Sigma^{\prime}$ and multiplied with its opacity value $\sigma$. 3D Gaussian parameters $\mu,q,S,\sigma,\xi$ are optimized during training, and the number of 3D Gaussians is adaptively controlled.

### v2

# Learning High-Fidelity Robot Self-Model with Articulated 3D Gaussian Splatting

@@ Abstract
Self-modeling enables robots to learn task-agnostic models of their morphology and kinematics based on data that can be automatically collected, with minimal human intervention and prior information, thereby enhancing machine intelligence. Recent research has highlighted the potential of data-driven technology in modeling the morphology and kinematics of robots. However, existing self-modeling methods suffer from either low modeling quality or excessive data acquisition costs in equipment. Beyond morphology and kinematics, surface color is also a crucial component of robots, which is challenging to model and remains unexplored. In this work, a high-quality, surface color-aware, and link-level method is proposed for robot self-modeling. We utilize three-dimensional (3D) Gaussians to represent the static morphology and surface color of robots, and cluster the 3D Gaussians to construct neural ellipsoid bones, whose deformations are controlled by the transformation matrices generated by a kinematic neural network. The 3D Gaussians and kinematic neural network are trained using data pairs composed of joint angles, camera parameters, and multi-view images without depth information. By feeding the kinematic neural network with joint angles, we can utilize the well-trained model to describe the corresponding morphology, kinematics, and surface color of robots at the link level, and render robot images from different perspectives with the aid of 3D Gaussian splatting. Furthermore, we demonstrate that the established model can be exploited to perform downstream tasks such as motion planning and inverse kinematics.
@@ 1 Introduction
High-quality models of geometry and kinematics are essential for effective robot planning, control, and visualization. Traditional modeling approaches demand substantial human effort to gather the necessary parameter information of the robot, and over time, the quality of these models tends to deteriorate due to various factors, such as unexpected damage, joint malfunctions, and wear. Therefore, the development of methods that enable robots to collect data and learn their own models autonomously is critical for ensuring the successful execution of tasks and advancing the level of machine intelligence. Inspired by human behavior (Rochat, 2003), the concept of robot self-modeling has recently gained attention among researchers (Chen et al., 2022; Díaz Ledezma and Haddadin, 2023). A robot self-model is a general-purpose representation of a robot’s physical shape and structure that can be acquired at any time without human intervention (Kwiatkowski, 2022; Schulze and Lipson, 2024). Through self-modeling, robots can autonomously and continuously update their knowledge of both morphology and kinematics, which not only ensures that the model information remains up-to-date and adaptable to various unforeseen circumstances but also simplifies the modeling process for newly designed robots.
@@
Early approaches focused on designing task-specific self-modeling techniques, (Chen et al., 2021; Hang et al., 2021) such as modeling end-effector positions (Mathew et al., 2014), tilt angles (Bongard et al., 2006), or joint velocities (Sanchez-Gonzalez et al., 2018) according to task requirements. However, due to the lack of generality and adaptability in these methods, recent work has shifted toward modeling the whole robot to construct task-agnostic self-models (Kwiatkowski and Lipson, 2019; Chen et al., 2022). They employed visual data with depth information to enable self-modeling of the overall morphology of a 4-degree-of-freedom robot.  Yang et al. (2024a) later adapted it to robots with higher degrees of freedom. Chen et al. (2022)’s approach allows the robot to determine its spatial occupancy based on input joint configurations but cannot directly compute the kinematic information (e.g., position) of specific links. To address link-level modeling,  Yang et al. (2024a) developed multiple independent models for each robot link and integrated them into a comprehensive self-model. However, these methods require depth information to create self-models, which necessitates devices with higher cost compared to standard RGB cameras (e.g., depth cameras and LiDAR), and is hard to obtain in practice. In addition to visual data, inertial measurement unit (IMU) data can also be utilized for the self-modeling of robots (Díaz Ledezma and Haddadin, 2023), which necessitates the attachment of numerous IMU sensors to robot links. Notably, despite the considerable advancements in robot self-modeling concerning morphology and kinematics, the aforementioned approaches have neglected the modeling of surface color, which is crucial for the visualization of robots in simulation environments and digital twin systems. When visualizing simulation results, a robot’s 3D model with surface color is more conducive to understanding and more human-friendly. Currently, commonly used simulators are typically equipped with robot models with surface color, such as MuJoCo(Todorov et al., 2012), Gazebo(Koenig and Howard, 2004), and Nvidia Isaac Sim(Liang et al., 2018). Thus, beyond morphology and kinematics, surface color modeling remains a significant challenge in achieving comprehensive robot self-modeling.
@@
Given the challenges and high costs in equipment for data acquisition in the previously discussed self-modeling methods and the importance of surface color modeling, a visual self-modeling approach utilizing only two-dimensional (2D) images presents a more efficient and cost-effective alternative. However, the current 2D image-based self-modeling method suffers from poor morphology modeling quality and fails to model the link-based structure of robots (Schulze and Lipson, 2024). In this paper, we develop a scheme based on 3D Gaussian Splatting (3DGS) (Kerbl et al., 2023) and neural networks for modeling the robot’s morphology, kinematics, and surface color. 3DGS is a technology for 3D scene reconstruction, which represents 3D scenes with a large number of optimizable, flexible, and expressive 3D Gaussian ellipsoids. Recent improvements equip it with the ability to real-time render scenes into images (Kerbl et al., 2023). With the assistance of 3DGS, we can acquire 3D information, which is necessary for robot self-modeling from 2D images. In order to control 3D Gaussian ellipsoids to reflect the robot pose, we use a kinematic network to move them. An overview of our method is demonstrated in Figure 1(a). To optimize the 3D Gaussians and kinematic network, the robot only needs to gather multi-view RGB images under various configurations, along with corresponding joint configuration data and camera parameters. The static morphology and surface color information of the robot’s links are embedded in 3D Gaussians, each of which is an ellipsoid defined by its position, covariance matrix, color, and opacity. By clustering these 3D Gaussians, the bones of the robot model are constructed, with transformations of the bones determined by the kinematic network. By feeding the kinematic neural network with joint angles, we can utilize the well-trained self-model to describe the corresponding morphology, kinematics, and surface color of robots at the link level and render robot images from different perspectives with the aid of 3DGS. Besides, the self-model can be employed to directly control the robot to perform downstream tasks such as motion planning and inverse kinematics. As illustrated in Figure 1(b), compared to previous methods, the proposed method produces a high-quality and link-level self-model of robots.
@@
Figure 1: Robot self-modeling from images. a Given a robot whose model is unknown, our method can build a self-model of the robot from multiple RGB images. We capture multi-view photos of the robot at different joint configurations, and train the 3D Gaussians and the kinematic network, which are responsible for representing the static morphology and surface color of robots and neural bone deformation, respectively. The well-trained model can be utilized to render robot images with the aid of 3D Gaussian splatting and perform downstream tasks. b Compared to the methods proposed by Chen et al. (2022) and Schulze and Lipson (2024), our method can learn a high-quality, surface color-aware, and link-level self-model.
Overall, the contributions of this paper are summarized as follows.
  • A novel robot self-modeling method based on 3D Gaussian Splatting is proposed, which enables efficient reconstruction from multi-view RGB images. This is one of the first methods leveraging 3DGS in robot self-modeling.
  • The robot can be modeled at the link level. With the predefined link number of the robot, the self-model automatically identifies the spatial positions and kinematic connections of these links based on the learnable and controllable neural bone architecture. This is the first method to use such architecture in robot self-modeling.
  • Our proposed method achieves state-of-the-art performance, outperforming previous works of robot self-modeling in terms of geometric accuracy.
@@
  • The self-model can be utilized to control the robot. Experiments show that the self-model can be utilized to control the robot to accomplish downstream tasks without additional training.
@@ 2 Related work
Neural Radiance Field (NeRF) has shown remarkable results in reconstructing scenes from multiple-view photos. It uses an implicit function and volume rendering to represent a static scene. Spatial coordinates $x$ and view directions $d$ are mapped into corresponding density $\sigma$ and color $c$. All points along a ray that connects a pixel on the image and the camera’s optical center are rendered into a pixel on the image plane by volume rendering, which is formulated as
$C(r)=\sum_{i=1}^{N}\exp\left(-\sum_{j<i}\sigma_{j}^{r}\delta_{j}^{r}\right)\left(1-\exp\left(-\sigma_{i}^{r}\delta_{i}^{r}\right)\right)c_{i}^{r}$ (1)
where $\delta_{i}^{r}$ is the distance between two sampled points along the ray, and $C(r)$ is the color on the rendered image. However, during both training and rendering, it is necessary to sample hundreds or even thousands of points along each ray and perform a time-consuming network query for every single point to obtain its color and density, resulting in an extremely high computational cost. As the original NeRF is for static scenes only, Pumarola et al. (2021) proposed the first method to model dynamic scenes with NeRF. The key idea is to bend the ray and map the moved points to points on the original ray.
Some methods have been proposed to accelerate the training of NeRF (Fang et al., 2022; Müller et al., 2022). These works motivate researchers to use dynamic NeRF to reconstruct and animate the human body (Noguchi et al., 2021) and animals (Yang et al., 2022). Noguchi et al. (2022) combined NeRF with signed-distance functions (SDFs) to learn both the appearance and the structure of articulated objects, including robots and human bodies, by observing their movements from multiple views. They modeled articulated objects with several ellipsoids, and could repose articulated objects by transforming ellipsoids. Uzolas et al. (2023) extracted point clouds of articulated objects from pre-trained NeRF models and used Linear Blend Skinning (LBS) to learn dynamic NeRF and associated skeletal models jointly. Although these works achieved modeling robots from images or videos, they did not associate the robots’ joints with the built morphology models. In other words, they did not directly use joint angles as the primary deformation inputs to train their model and thus failed to model the robot kinematics described by joint angles. Controlling the model learned by their method requires more information, such as link length, link twist, and joint offset, in addition to joint angles.
@@
Though NeRF shows promising results in 3D reconstruction, it has a significant drawback. Training and rendering a high-quality scene reconstruction may take a long time. The 3DGS method is designed to solve this problem and can quickly complete training and learn high-quality models to present well-rendered images with a high peak signal-to-noise ratio. Similar to NeRF, researchers extended 3DGS to the reconstruction of dynamic 3D scenes (Wu et al., 2024; Yang et al., 2024c; Liang et al., 2025; Yang et al., 2024b; Lin et al., 2024). For example, Hu et al. (2024) and Qian et al. (2024) proposed to model 3D human avatars with 3DGS. These works require a Skinned Multi-Person Linear (SMPL) model as a base human model (Loper et al., 2023). Several other approaches aim to design controllable and editable models with 3DGS. Yu et al. (2024) used predefined masks to choose moving parts of the modeled scene. These masks are used to control the model, and only the chosen part can be controlled. Huang et al. (2024) used control points to control the dynamic 3DGS model. Control points are automatically generated during training, which then control the movement of the whole 3DGS model with Linear Blend Skinning(LBS) (Sumner et al., 2007). Zhang et al. (2024) proposed to use neural bones to control the 3DGS model, which was similar to the method proposed by Yang et al. (2022). A neural bone is an ellipsoid parameterized by position and radius. Nevertheless, these works still did not use the joint values between links of an articulated object as inputs to the model, and none of these works is designed for the purpose of robot self-modeling.
2.3 Robot self-modeling with 2D input
@@
The methods mentioned above are primarily designed to model humans and animals, and are not intended for robot self-modeling. Several studies have explored using 2D images for robot self-modeling (Hu et al., 2023; Marques Monteiro et al., 2024). However, these approaches are either only applicable to simple robots with few degrees of freedom or limited to rendering images of the robot in various states, and do not succeed in constructing a comprehensive model of the robot’s morphology. Schulze and Lipson (2024) was the first to solve the robot self-modeling problem using dynamic NeRF and associated the built model with input robot joints. However, NeRF-based self-modeling requires optimizing a dense multi-layer perceptron (MLP) for learning both morphology and kinematics, and uses volumetric ray-marching to render robot images. The MLP expression makes the learned model difficult to be directly explained and edited. The ray-based volume rendering is also time-consuming. As a result, NeRF-based self-modeling suffers from problems such as being difficult to balance training time and modeling quality. Another limitation of their approach is that they did not present results including estimated surface color, although the NeRF method they adopted is capable of rendering surface color. Lou et al. (2024) used a hybrid representation that combines 3DGS and mesh to create digital assets of robots. The model of robots can be controlled by the robot joints. However, real Denavit-Hartenberg parameters are required to control the model based on joints in training and inference stages, which requires human intervention and fails to model robot kinematics automatically.
A concurrent work (Liu et al., 2024) also achieved modeling robots with RGB images only and maps joint values to the movement of the model. However, their method requires a known connection order of the robot’s links to calculate the forward kinematics. Our method only requires the number of robot links.
In addition, Labbé et al. (2021) and Lu et al. (2023) also involve shape reconstruction and differentiable rendering. They focus on estimating the unknown joint angles of a robot given a known robot model and images of the robot. However, their work did not consider an unknown robot, which represents a fundamental difference from our approach. Our work focuses on learning an unknown model of a robot based on its images with joint annotations. A geometric model of the robot is also learned to visualize the learned robot’s kinematics. While both papers and our work render images of robots, the purposes are different. Specifically, the rendered images in their work are generated based on estimated joint angles and the known robot model. The primary purpose of comparing these rendered images with real images is to quantify the accuracy of the estimated joint angles. In our work, image rendering is performed based on known joint angles and the learned geometric and kinematic models, for the purpose of evaluating the accuracy of the learned models. Even though Lu et al. (2023) learned a mesh reconstruction, they did not focus on the accuracy of the learned mesh reconstruction compared with a ground-truth robot mesh. Therefore, directly comparing the quality of rendered images of our work and theirs does not make sense because tasks and prior conditions are totally different.
@@ 3 Method
Figure 2: Overview of our method. We first train a static 3DGS self-model with images of the robot at zero pose from various views. Then, we initialize neural ellipsoid bones to model the robot kinematic chain. Using images of the robot at different joint configurations, we train a dynamic self-model. The neural bones are deformed, and the relative 3D Gaussian is controlled by LBS. Finally, the model is rendered into images and compared with the ground truth images.
We propose a novel robot self-modeling method that begins by modeling the robot in a rest pose, which we define as the canonical space. The robot is then transformed into various poses that form the transformed space. To implicitly represent the self-model of the robot, we utilize 3DGS. The model can be deformed and rendered into images, allowing us to visualize the robot’s morphology from different perspectives. An overview of our method is provided in Figure 2.
@@ 3.1 Preliminaries
3DGS uses 3D Gaussians to represent a static scene (Kerbl et al., 2023). A 3D Gaussian is a small ellipsoid with feature parameters. Typically, a static 3DGS model consists of 10 thousand 3D Gaussians. Each 3D Gaussian is defined by a full 3D covariance matrix $\Sigma$:
$G(x)\ =\ e^{-\frac{1}{2}{x}^{T}\Sigma^{-1}x}$ (2)
where $x\in\mathbb{R}^{3}$ denotes the position coordinate of a point in the ellipsoid. The covariance matrix $\Sigma$ can be calculated using the rotation matrix $R$ and the scaling matrix $S$ as $\Sigma=RSS^{T}R^{T}$. The rotation matrix is derived from a quaternion $q$. Each Gaussian has an opacity value $\sigma$ and a spherical harmonic coefficient $\xi$ to create a view-dependent color. When rendering images, 3D Gaussians are projected to a 2D plane. Given a viewing transformation $W$, the covariance matrix $\Sigma^{\prime}$ in camera coordinates is given by
@@
where $J$ is the Jacobian of the affine approximation of the projective transformation. The color $C$ of a pixel is computed by alpha blending 3D Gaussians that overlap on the pixel sorted by depth, which is given by
@@
where $N$ is the number of Gaussians overlapping the pixel, $c_{i}$ is the color of Gaussian $i$ calculated by spherical harmonic coefficient, $\alpha$ is given by evaluating the projected Gaussian with covariance matrix $\Sigma^{\prime}$ and multiplied with its opacity value $\sigma$. 3D Gaussian parameters $\mu,q,S,\sigma,\xi$ are optimized during training, and the number of 3D Gaussians is adaptively controlled. By representing scenes with a direct and explicit model of millions of 3D Gaussians, it can be rendered efficiently in a single pass via a splatting method—analogous to traditional graphics rasterization and highly optimized for modern GPUs—thereby achieving both training and rendering speeds.